### v2

# Induced Gravitational Waves as Cosmic Tracers of Leptogenesis

Abstract
We demonstrate that induced gravitational waves (IGWs) can naturally emerge within the framework of thermal leptogenesis models, thereby providing a robust probe for exploring this theory at remarkably high energy scales. To illustrate this principle, we put forth a basic leptogenesis model in which an early matter-dominated phase, tracing the leptogenesis scale, enhances the generation of gravitational waves induced by an early structure formation. Leveraging recent N-body and lattice simulation results for IGW computations in the non-linear regime, we show that it is possible to establish a direct link between the frequency and amplitude of these IGWs and the thermal leptogenesis scale.
Introduction.  Thermal leptogenesis is one of the most studied mechanisms for generating the Baryon Asymmetry of the Universe (BAU), owing to its simplicity and its direct link to low-energy neutrino physics [1, 2, 3]. While neutrino observables provide valuable insights into thermal leptogenesis given a proper treatment of the theory’s flavor structure [4, 5, 6, 7], a laboratory probe of the leptogenesis scale remains infeasible. This is because, in its simplest form, thermal leptogenesis occurs at extremely high temperatures, namely $T_{\text{lepto}}\sim M_{N}\gg\rm TeV$, where $M_{N}$ is the mass scale of the heavy right-handed neutrino (RHN), introduced on top of the Standard Model (SM) to generate light neutrino masses and facilitate BAU. Notably, we may probe such high energy scales with gravitational waves (GWs), thanks to the remarkable theoretical and experimental advancements in detecting the stochastic GW background [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. As such, the LIGO-Virgo-KAGRA (LVK) collaboration has set upper bounds on the stochastic gravitational wave background at $\mathcal{O}(\text{Hz})$ frequencies [8]. Recent Pulsar Timing Array (PTA) observations indicate a nHz stochastic signal that may be of primordial origin [14, 15, 16, 17, 18]. Mid-frequency detectors like LISA, scheduled for launch in the 2030s, will feature a comprehensive physics pipeline [19, 20, 21, 22, 23].
In this Letter, we investigate the production of GWs in thermal leptogenesis models and argue that their frequency spectrum may encode information about the thermal leptogenesis scale. In particular, we focus on the induced gravitational waves (IGWs), sourced inevitably by primordial fluctuations in the very early universe due to second order gravitational interactions [24, 25, 26, 27, 28, 29, 30] (see Ref. [31] for a review). Interestingly enough, the presence of an early matter-dominated (eMD) epoch can intensify the production of IGWs [32, 30, 33, 34, 35, 36, 37, 38, 39, 40, 41, 31, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. As we demonstrate here, simple leptogenesis models possess all the necessary prerequisites to achieve such an eMD era well-before Big Bang Nucleosynthesis (BBN), that is at $T_{\rm BBN}\gtrsim 4\,\text{MeV}$ [53, 54, 55, 56, 57], making IGWs a powerful probe of thermal leptogenesis.
Accurate estimations of the IGW spectrum generated during an eMD epoch are challenging; density fluctuations grow, and non-linear structures may emerge. Moreover, the specifics of transition to the following radiation-dominated era can impact the final IGW spectrum [37, 38, 50]. One may restrict calculations to scales within the linear regime [36, 37, 38], where analytical or semi-analytical estimates are reliable. However, one expects a louder, more interesting GW signal in the non-linear regime [33, 34], though accurate predictions require numerical simulations [43, 48, 52]. Recently, Ref. [48] performed hybrid N-body and lattice simulations, following the formation and decay of early structures, and computed the resulting IGW spectrum.
Focusing in this work on the non-linear regime, we chart the parameter space of simple thermal leptogenesis models that yield a long enough eMD epoch with a detectable IGW signal. Furthermore, we argue that there is a concrete link between the leptogenesis scale and the IGW spectrum. Thus, a detection of an IGW background would hint at a particular leptogenesis scale. In any event, our work will serve to exclude a significant portion of our leptogenesis model parameter space.
Early matter-dominated epoch in thermal leptogenesis models.  We consider the widely studied ultraviolet realization of seesaw models based on the gauge $U(1)_{B-L}$ symmetry with coupling $g^{\prime}$, naturally embedded in many Grand Unified Theories (GUTs) [58, 59, 60]. The relevant terms of the Lagrangian read as
$-\Delta\mathcal{L}$ $\subset$ $\frac{1}{2}y_{N}\overline{N_{R}}\Phi N_{R}^{C}+y_{D}\overline{N_{R}}\tilde{H}^{\dagger}L$ (1)
$+\frac{1}{4}\lambda_{H\Phi}H^{2}\Phi^{2}+V(\Phi,T)\,$,
where the SM lepton doublets $L$, the SM Higgs doublet $H$, the right-handed neutrino fields $N_{R}$ and the scalar field $\Phi$ have BLB-L charges -1, 0, -1 and 2, respectively. The first two terms are relevant for the masses of light and heavy neutrinos. After the $U(1)_{B-L}$ phase transition at a critical temperature $T_{c}$ (discussed later), $\Phi$ gets its vacuum expectation value $v_{\phi}$, and RHNs become massive. For simplicity, we assume a single RHN mass scale $M_{N}=y_{N}v_{\phi}$ with $v_{\Phi}=\mu/\sqrt{\lambda}$ determined from the zero-temperature potential $V(\Phi,0)=-\mu^{2}\Phi^{2}/2+\lambda\Phi^{4}/4$. These massive RHNs then decay CP-asymmetrically into lepton doublets and Higgs bosons, producing the lepton asymmetry around the temperature $T_{\rm lepto}\sim M_{N}$–owing to the chosen one-scale leptogenesis model, the final BAU needs to be computed within the resonant leptogenesis framework with quasi-degenerate RHNs [61]. Later on, once the SM Higgs takes its vacuum expectation value $v_{h}=174~{\rm GeV}$ at the electroweak phase transition, light neutrino masses $m_{\nu}\sim{y_{D}}^{2}v_{h}^{2}/M_{N}$ are generated via the type-I seesaw mechanism [62, 63, 64, 65]. The third term in Eq. (1) is the Higgs portal interaction, responsible for the decay $\Phi\to hh$, which may occur before or after the electroweak phase transition. Finally, the last term is the finite-temperature potential describing the $\Phi$ dynamics and the breaking of the $U(1)_{B-L}$ symmetry [66, 67, 68, 69, 70].
At very high temperatures, the $U(1)_{B-L}$ symmetry is unbroken and the potential has a minimum at $\Phi=0$. At lower temperatures, a secondary minimum is created at $\Phi\neq 0$, causing a potential barrier. The two minima become degenerate at the critical temperature $T_{c}$ and, for $T\lesssim T_{c}$, the barrier height decreases and disappears at $T_{*}$, making $\Phi=0$ into a maximum. The scalar field then transitions to $v_{\Phi}$. The transition strength is roughly characterized by the order parameter $\Phi_{c}/T_{c}$ with $\Phi_{c}=\Phi(T_{c})$ [68]. For $\Phi_{c}/T_{c}\ll 1$, the barrier vanishes quickly $(T_{c}\simeq T_{*}=2\sqrt{g^{\prime}}v_{\Phi})$, allowing for smooth rolling from $\Phi=0$ to $\Phi=v_{\Phi}$. We consider values of $\lambda$ and $g^{\prime}$ that satisfy this condition, specifically $\lambda\simeq{g^{\prime}}^{3}$ and $g^{\prime}\lesssim 10^{-1}$. This relation serves as a useful rule of thumb: increasing the power of $g^{\prime}$ raises the barrier height, hindering smooth rolling, while lowering it is disfavored by constraints discussed later. However, variations around $\lambda\sim{g^{\prime}}^{3}$ do not qualitatively alter the model, as the most sensitive parameter is the coupling $y_{N}$.
At temperatures $T<T_{*}$, the scalar field undergoes coherent oscillations around $v_{\Phi}$ with an angular frequency $m_{\Phi}=\sqrt{2\lambda}v_{\Phi}$, behaving like a non-relativistic matter component [71, 72]. If these oscillations persist due to a sufficiently long lifetime of the scalar field, the Universe undergoes an eMD epoch which starts and ends at the temperatures $T_{\rm dom}$ and $T_{\rm dec}$, respectively. The former is given by $T_{\rm dom}=(\rho_{\Phi}(T_{c})/\rho_{R}(T_{c}))T_{c}$ with $\rho_{\Phi(R)}$ denoting the field (radiation) energy density, while the latter depends on the scalar field decay.
Figure 1: The green shaded region shows the parameter space linking the eMD epoch to the leptogenesis scale in the plane $M_{N}$-$y_{N}$ for $g^{\prime}=10^{-2}$. The black lines bound the allowed parameter space according to the constraints described in the main text. The color gradient displays the minimal value of the scalar amplitude required to produce IGWs (see Eq. (4)). The top x-axis reports the peak GW frequency $f_{\rm high}$ defined in Eq. (7) computed for $A_{\rm s}=A_{\rm s}^{\rm min}$. The points highlight the benchmark cases in Fig. 2.
In our scenario, the decay rate of $\Phi$ is governed by the Higgs portal coupling $\lambda_{H\Phi}\simeq\lambda_{H\Phi}^{\rm tree}+\lambda_{H\Phi}^{\text{1-loop}}$, which receives contributions from both tree-level and the inevitable one-loop effects. Note that $\lambda_{H\Phi}^{\text{1-loop}}$ is the minimal value of the Higgs portal coupling in seesaw models described by the first two terms in Eq. (1) [73, 74, 75]. Most interesting for our purposes is the case where $\lambda_{H\Phi}^{\text{1-loop}}>\lambda_{H\Phi}^{\rm tree}$ as it establishes a direct link between the leptogenesis scale (and so the neutrino parameters) and the lifetime of $\Phi$—thereby the IGW spectrum.
To ensure such a connection, we consider other decay channels negligible. The decays $\Phi\to Z^{\prime}Z^{\prime}$ and $\Phi\to NN$ are kinematically forbidden by requiring $M_{Z^{\prime}}=\sqrt{2}g^{\prime}v_{\Phi}>m_{\Phi}$ and $M_{N}\gtrsim m_{\Phi}$, respectively. The latter condition also rules out any off-shell, one-loop processes involving $N$ or $Z^{\prime}$ in the final state. Additionally, the one-loop $\Phi\Phi ZZ$ coupling, arising from the kinetic mixing [76, 77, 78, 79], is much weaker to produce any considerable effect. Another competitive channel in this model is the three-body decay to SM fermions and vector bosons, $\Phi\to f\bar{f}V$, mediated by virtual $Z^{\prime}$ at one-loop [80, 81].
In the limit $m_{\Phi}\gg m_{h}$, the one-loop $\Phi\to hh$ decay rate is given by [73, 74, 75, 82, 83]
$\Gamma_{\Phi}^{hh}\simeq 1.8~{\rm MeV}~\frac{y_{N}^{3}}{\lambda^{1/2}}\left(\frac{M_{N}}{10^{11}~\text{GeV}}\right)^{3}~\ln^{2}\left(\frac{\Lambda_{\rm GUT}}{\Lambda}\right)\,$, (2)
where the logarithmic dependence accounts for the renormalization condition from the GUT scale, $\Lambda_{\rm GUT}\simeq 10^{16}~\text{GeV}$, to a lower scale $\Lambda$ [73, 74]. Note that while deriving Eq. (2), we use the seesaw relation $y_{D}=\sqrt{m_{\nu}M_{N}/v_{h}^{2}}$, with $m_{\nu}\simeq 0.01$ eV. This implies that the lifetime of $\Phi$ also depends on the light neutrino masses, making the model strongly connected to neutrino oscillation data and the upper bound $\sum m_{\nu}\lesssim 0.2~{\rm eV}$ [84, 85]. The decay temperature $T_{\text{dec}}$ can be determined by implicitly solving ${\rm H}(T_{\rm dec})\simeq\Gamma_{\Phi}^{hh}(T_{\rm dec})$, where $\rm H$} is the Hubble parameter and $\Lambda=T_{\rm dec}$.
In Fig. 1, we show without loss of generality the parameter space for $g^{\prime}=10^{-2}$ achieving an eMD era uniquely determined by the two remaining model parameters, i.e., the leptogenesis scale $M_{N}$ and the coupling $y_{N}$. We later discuss variations of $g^{\prime}$ (see also the Supplemental Material). Several necessary conditions tightly constrain the target parameter space.
First, the scalar field must dominate the Universe for long enough before decaying in order to develop non-linear structures. This requirement is related to the amplitude of the primordial spectrum of density fluctuations, say $A_{\rm s}$. Thus, given a duration of the eMD era, we may set a minimum value for $A_{\rm s}$ (which we define as $A^{\rm min}_{\rm s}$ in Eq. (4)) and vice-versa. We take $A^{\rm min}_{\rm s}\leq 10^{-2}$ as a generous upper bound not to overproduce Primordial Black Holes (PBHs), see e.g., Refs. [86, 87, 88, 89] for a general review.
Second, the coupling $y_{N}$ is bounded from below and above by demanding $m_{\Phi}\gtrsim M_{N}$ (prohibiting $\Phi\to NN$) and $T_{\rm lepto}\lesssim T_{c}$ (allowing the production of the lepton asymmetry), respectively. Laslty, the decay channel $\Phi\rightarrow hh$ must dominate over $\Phi\rightarrow f\bar{f}V$ ($\Gamma_{\Phi}^{hh}\gtrsim\Gamma_{\Phi}^{f\bar{f}V}$). Other constraints such as $T_{\rm dec}\gtrsim T_{\rm BBN}\sim 4~\text{MeV}$ and $v^{\rm max}_{\Phi}\simeq 5\times 10^{15}~\text{GeV}\lesssim\Lambda_{\rm GUT}$ with the scale separation justified by the seesaw-perturbativity condition [90, 91, 92], are less restrictive for $g^{\prime}=10^{-2}$. Let us now move on to the GW signal.
Induced gravitational waves.  Bulk velocities in the Universe lead to anisotropic stresses which then source GWs (since $T_{ij}\sim\rho v_{i}v_{j}$ where $v_{i}$ is the spatial velocity). During an eMD phase, density fluctuations and velocity flows grow, especially in the non-linear regime where structures form (here, $\Phi$-halos). The scale of non-linearities at a given conformal time $\tau$ is estimated by [32, 36]
$k_{\rm NL}(\tau)\sim\alpha\,A_{\rm s}^{-1/4}{\cal H}\,$, (3)
where we assumed a scale-invariant spectrum of curvature fluctuations with amplitude $A_{\rm s}$ and $\alpha\approx 1.7$, though the exact value of $\alpha$ depends on whether one uses the Poisson equation [36] or spherical collapse criterion [48]. Fourier modes with $k>k_{\rm NL}$ are in the non-linear regime.
In the linear regime, the IGW spectrum produced is roughly given by $\Omega_{\rm GW}\sim A_{\rm s}^{2}$ for $k<k_{\rm NL}$, if the transition to radiation-domination takes at least one $e$-fold [38], as in our case. The observationally relevant regime then corresponds to $A_{\rm s}\sim 10^{-1}-10^{-5}$. However, Eq. (3) implies that the smallest scale that becomes non-linear at the time of $\Phi$-decay is $k_{\rm NL}(\tau_{\rm dec})/{\cal H}_{\rm dec}\sim 3-30$. Thus, modes in the linear regime with potentially observable GWs do not experience much of the eMD phase.
In the non-linear regime, as shown in Ref. [48], the IGW generation is dominated by the largest structures that form at the end of the eMD era. This results in an amplitude of the GW spectrum which is largely independent of the reheating timescale [48], provided that the eMD phase lasts long enough. Concretely, we need at least that the largest possible non-linear scale enters the Hubble radius during the eMD, namely $k_{\rm NL}(\tau_{\rm dec})<{\cal H}_{\rm dom}$. Using thus that $a\sim\tau^{2}$ in the eMD epoch, the aforementioned inequality translates into a lower bound on $A_{\rm s}$ given by [48]
$A_{\rm s}^{\rm min}\equiv\alpha^{4}\left(\frac{a_{\mathrm{dom}}}{a_{\mathrm{dec}}}\right)^{2}\approx 9\left(\frac{T_{\mathrm{dec}}}{T_{\mathrm{dom}}}\right)^{2}\,.$ (4)
Since the analytical treatment in the non-linear regime is challenging, we rely on the results of numerical simulations [43, 48, 52] for precise characterization of the GW spectrum. In particular, we use the results of Ref. [48], which include the state-of-the-art fit to the GW spectrum induced by a scale-invariant spectrum at the end of the eMD epoch, namely
$\Omega_{\mathrm{GW}}(k)\simeq 0.05\,A_{\mathrm{s}}^{7/4}\left(\frac{k}{\mathcal{H}_{\mathrm{dec}}}\right)^{3/2}\,.$ (5)
This spectrum has low- and high-$k$ cut-offs given by $k_{\rm low}\approx 15{\cal H}_{\rm dec}$ and $k_{\rm high}\approx 9k_{\rm NL}(\tau_{\rm dec})\approx 14\mathcal{H}_{\mathrm{dec}}/A_{\mathrm{s}}^{1/4}$, respectively. The former is due to numerical resolution limitations [48] while the latter is due to possible artefacts from coherent effects of non-relativistic $N$-body simulations and given by the light-crossing time of the largest halos. Generally, though, one expects a smooth decay of the GW spectrum rather than a sharp cut-off at $k_{\rm high}$. Refs. [41, 44, 43, 48, 52] suggest a high-frequency power law decay as $\Omega_{\mathrm{GW}}\propto 1/f^{n}$ with $n\gtrsim 1$, though fully relativistic simulations are necessary to confirm this [93]. The results of linear perturbation theory are recovered at low enough frequencies. It should be noted that the $A_{\mathrm{s}}$-dependence of the GW spectrum in Eq. (5) agrees with the Press-Schechter argument of halo collapse of Ref. [52].
From Eq. (5) we see that once we know the peak amplitude of the GW spectrum, which only depends on $A_{\rm s}$, we can tell apart the value of ${\cal H}_{\rm dec}$, and so $T_{\rm dec}$, from the peak frequency. This provides a remarkable link to leptogenesis, since in our scenario $T_{\rm dec}$ is related to $\Gamma_{\Phi}^{hh}$ via Eq. (2). For the values of the leptogenesis parameters compatible with such a GW signal see Fig. 1.
For the reader’s convenience, we express the GW spectrum today and frequencies in terms of typical parameters reconstructible within our leptogenesis model, namely
$\Omega_{\rm GW,0}h^{2}=4.2\times 10^{-5}\,A_{s}^{11/8}\left(\frac{f}{f_{\mathrm{high}}}\right)^{3/2}\,$, (6)
where we account for the redshifting factor assuming the SM effective degrees of freedom,
$f_{\rm high}\simeq 6.4\times 10^{-5}\text{ Hz}\left(\frac{A_{\rm s}}{10^{-5}}\right)^{-1/4}\left(\frac{T_{\mathrm{dec}}}{10\,\text{GeV}}\right)\,$, (7)
$f_{\rm low}\simeq 3.8\times 10^{-6}\text{ Hz}\left(\frac{T_{\mathrm{dec}}}{10\,\text{GeV}}\right)\,.$ (8)
We note from Eq. (6) that the CMB normalization for primordial fluctuations, i.e. $A_{\mathrm{s}}\sim 10^{-9}$ [94], yields $\Omega_{\rm GW,0}h^{2}\sim 10^{-17}$. Thus, a detectable IGW signal indicates enhancement of fluctuations during inflation (see Ref. [95] for a review).
Results and discussion.  In our model, the minimal Higgs portal coupling predicts a link between the leptogenesis scale $M_{N}$ and the IGW spectrum, as it determines the end of the eMD epoch (cf. Eq.(2)) and the peak GW frequency. Namely, by observing the peak of the IGW spectrum, we can deduce $T_{\rm dec}$ from the peak frequency and so $M_{N}$. The duration of the eMD is jointly determined with $y_{N}$, see Fig. 1, further shrinking the target parameter space.
In Fig. 2 we show the GW spectrum predicted by our leptogenesis model for different benchmark scenarios. The solid lines display Eq. (6) between the cut-off frequencies $f_{\rm low}$ and $f_{\rm high}$, while the dashed lines show the continuation $\Omega_{\rm GW,0}(f>f_{\rm high})\propto f^{-1}$ [44, 52] (see discussion below Eq. (5)). The first four benchmark points correspond to scenarios with $g^{\prime}=10^{-2}$ and $A_{\rm s}=A_{\rm s}^{\rm min}$ (see Fig. 1). Hence, they represent the most conservative scenarios in terms of detectability, as values $A_{\rm s}>A_{\rm s}^{\rm min}$ would imply higher GW amplitudes. Specifically, BP1 represents the lowest allowed GW signature for $g^{\prime}=10^{-2}$ ($A_{\rm s}^{\rm min}\simeq 10^{-5}$), while BP2 and BP3 correspond to the highest GW signatures having $A_{\rm s}^{\rm min}\simeq 10^{-2}$. Notably, for realistic models with large gauge couplings, $A_{\rm s}\sim 10^{-5}$ serves as an absolute lower bound (see Fig. 1).
We find that well-separated leptogenesis scales lead to significantly different $T_{\rm dec}$ values and, consequently, distinct peak frequencies $f_{\rm high}$, as also highlighted by the top x-axis in Fig. 1. Moreover, the smaller the GW energy density, the wider the range of validity of Eq. (6) with scaling as $f^{3/2}$. Thus, remarkably, we demonstrate that the measurement of the whole IGW spectrum would provide clear indications of the underlying leptogenesis/seesaw model.
As a compelling example of the extensive parameter space despite our model’s constraints, see the case BP5 in Fig. 2. By allowing for a smaller gauge coupling (e.g. $g^{\prime}=10^{-3}$ [96, 97, 98, 99]), our work suggests a plausible link between the GW background at the PTAs  [14, 15, 16, 17, 18] and leptogenesis models with $M_{N}\sim{\cal O}(10^{6}\,{\rm GeV})$, though a more precise estimate of the GW spectrum below the low frequency cut-off is required to draw definitive conclusions. This is possible because lower values for $g^{\prime}$ allow access to smaller decay temperatures $T_{\rm dec}$ and, consequently, a smaller frequency range, without violating the constraint from $\Phi\rightarrow f\bar{f}V$.
Figure 2: Benchmark IGW spectra from the eMD epoch driven by leptogenesis. The different lines corresponds to the benchmark points of the model parameter space reported in the top table. Except for BP5, they all assume $g^{\prime}=10^{-2}$ and $A_{\rm s}=A_{\rm s}^{\rm min}$ defined in Eq. (4). The full parameter space for $g^{\prime}=10^{-2}$ is provided in Fig. 1. The dashed lines display the extrapolation of the GW spectrum as $f^{-1}$. We also show the NANOGrav 15yr results [18] for the tentative nHz GW background [14, 15, 16, 17, 18]. The shaded regions represent the LVK bound [8] and the power-law integrated sensitivity curves for different experiments [100, 101].
In Fig. 3, we explore the correspondence between $M_{N}$ and $f_{\rm high}$ including $A_{\rm s}$. Specifically, the regions in different colors represent the allowed parameter space obtained by fixing the leptogenesis scale and varying the other parameters as $g^{\prime}\in[10^{-3.0},\,10^{-1.5}]$, $y_{N}\in[\sqrt{2g^{\prime 3}}\,,2\sqrt{g^{\prime}}]$ and $A_{\rm s}\in[\max(10^{-9},A_{\rm s}^{\rm min}),\,10^{-2}]$. We find that smaller $g^{\prime}$ values imply longer duration of the eMD and smaller $T_{\mathrm{dec}}$ without jeopardizing their parametric dependence on the $M_{N}$. This is evident from the enlarged regions corresponding to smaller $g^{\prime}$ values in Fig. 3. In this framework, leptogenesis scales $M_{N}\lesssim 10^{12}~\mathrm{GeV}$ have less discovery potential with IGWs owing to the constraint $v_{\phi}\leq v_{\phi}^{\rm max}$ and $A_{\rm s}\lesssim 10^{-2}$. We refer the reader to the Supplemental Material for further details, where we analyze the model parameter space corresponding to specific values of the signal amplitude and peak frequency, along with a discussion of potential theoretical uncertainties.
Let us also highlight two novel features of our leptogenesis framework. First, in this model, the leptogenesis scale $M_{N}$ is much higher than $T_{\rm dom}$. Therefore, in the computation of the BAU, one must account for the dilution factor $\Delta\simeq T_{\rm dec}/T_{\rm dom}$. The final value of the baryon-to-photon ratio in this case reads $\eta_{B}\simeq 10^{-3}\left(\frac{3m_{\nu}M_{Ni}}{8\pi v_{h}^{2}\delta}\right)\Delta$, where $\delta=(M_{Ni}-M_{Nj})/M_{Ni}$ accounts for the quasi-degeneracy among the RHNs [61, 82]. Requiring the observed value $\eta_{B}\simeq 6.3\times 10^{-10}$ and using Eq. (4), one finds $\delta\simeq 0.2\left(\frac{M_{Ni}}{10^{9}\,\mathrm{GeV}}\right)\sqrt{A_{\rm s}^{\rm min}}$. This implies that strong (weak) amplitude GWs are associated with weakly (strongly) quasi-degenerate RHNs. Second, since the GW frequency is sensitive to the leptogenesis scale, it follows that different flavor regimes, characterized by $M_{N}$ [4, 5, 6, 7], produce GWs at varying frequencies. This is significant because obtaining observable signatures through low-energy neutrino measurements typically requires imposing additional symmetries on the theory [102, 103, 104].
Figure 3: Allowed values for the GW amplitude in Eq. (6) computed at the peak frequency in Eq. (7), in case of different leptogenesis scales $M_{N}$ from $10^{6}~{\rm GeV}$ (left) to $10^{12}~{\rm GeV}$ (right). The color shading correspond to different intervals for the gauge coupling, with $A_{\rm s}\in[\max(10^{-9},A_{\rm s}^{\rm min}),\,10^{-2}]$. The black lines from left to right show the sensitivity of SKA, LISA, DECIGO, and ET detectors, respectively.
PBHs could be an additional signature of our scenario. Using the results of Refs. [86, 87] for the collapse of fluctuations in an eMD era, we see that amplitudes as low as $A_{\rm s}\sim 10^{-4}$ could lead to the formation of PBHs (see also the Supplemental Material), constituting potential viable candidates for dark matter. As the PBH fraction and mass depend on the duration and end of the eMD era, we leave a detailed study for future work. We caution though that there are still several uncertainties in PBH calculations, see e.g. Refs. [88, 105, 106], and that the PBH abundance is highly sensitive to primordial non-Gaussianities [107, 108, 109, 110].
In conclusion, with this Letter, we connected two seemingly unrelated phenomena: leptogenesis and IGWs. Our results go beyond the simple leptogenesis model considered here. Indeed, various seesaw mechanisms and leptogenesis models share similar Lagrangian constructions [111], likely containing long-lived fields. Thus, any beyond-the-Standard-Model physics framework involving an eMD epoch can be readily explored through Eqs. (4) and (6). Moreover, simple leptogenesis frameworks, as the one discussed here, might also produce cosmic strings that radiate GWs [90, 82], though a detailed study will be presented elsewhere.
It is remarkable that, assuming our leptogenesis model, a detection of such IGW background selects a preferred island in the parameter space and excludes all the rest. In the worst case scenario, even the absence of a GW signal constrains a significant region of the leptogenesis parameter space, inaccessible otherwise.
Acknowledgements.  We thank Satyabrata Datta and Gennaro Miele for insightful discussions. MC, RS, and NS acknowledge the support of the project TAsP (Theoretical Astroparticle Physics) funded by the Istituto Nazionale di Fisica Nucleare (INFN). The work of NS is further supported by the research grant number 2022E2J4RK “PANTHEON: Perspectives in Astroparticle and Neutrino THEory with Old and New Messengers” under the program PRIN 2022 funded by the Italian Ministero dell’Università e della Ricerca (MUR). GD is supported by the DFG under the Emmy-Noether program, project number 496592360, and by the JSPS KAKENHI grant No. JP24K00624. TP acknowledges the contribution of the LISA Cosmology Working Group as well as support of the INFN Sezione di Napoli through the project QGSKY.
Supplemental Material for
Induced Gravitational Waves as Cosmic Tracers of Leptogenesis
Marco Chianese, Guillem Domènech, Theodoros Papanikolaou, Rome Samanta, and Ninetta Saviano
This Supplementary Material extends and complements the discussion of the main Letter by first exploring more of the model parameter space with an induced GW signature, identifying degeneracies and possible theoretical uncertainties, and second discussing additional GW sources within our leptogenesis model. We do this separately below.
I Detailed parameter space exploration
I.1 Model parameter reconstruction and degeneracies
Our leptogenesis model is characterized by three parameters: $M_{N}$, $y_{N}$, and $g^{\prime}$. These determine both the onset and conclusion of the eMD phase at $T_{\rm dec}$ (and so its duration). The induced GW spectrum depends on the amplitude of primordial fluctuations $A_{s}$ and the reheating temperature $T_{\rm dec}$. Notably, the peak GW amplitude depends on $A_{s}$ only. Thus, given an observed GW spectrum, we can infer $A_{s}$ and $T_{\rm dec}$. We shall now illustrate the parameter space of the leptogenesis model that aligns with this scenario.
Before exploring the full parameter space, a few remarks are in order. Although some degeneracy is expected, since $T_{\rm dec}=T_{\rm dec}(M_{N},y_{N},g^{\prime})$, theoretical considerations significantly constrain the viable region. First, the length of the eMD phase must be long enough for non-linearities to develop, see Eq. (4). This means that the lower the $A_{s}$, the longer the eMD must be, and the smaller the allowed parameter space is. The parameter space closes at $A_{s}=A_{s}^{\rm min}$. Interestingly, when $A_{s}=A_{s}^{\rm min}(M_{N},y_{N},g^{\prime})$, the parameter space reduces to one dimension for a fixed GW spectrum. Namely, the model parameters $M_{N}$ and $y_{N}$ are fully specified for a given $g^{\prime}$. In other words, we obtain $M_{N}=M_{N}(g^{\prime})$ and $y_{N}=y_{N}(g^{\prime})$.
In Fig. S1, we show the allowed $M_{N}$$y_{N}$ parameter region after applying all model constraints, for $g^{\prime}=10^{-3}$ (left) and $g^{\prime}=10^{-4}$ (right). We note that the values of $y_{N}$ and $M_{N}$ are fully specified for a given set of $A_{s}^{\rm min}$ and $f_{\rm high}$. A notable feature is that decreasing $g^{\prime}$ opens the allowed parameter space to lower values of $y_{N}$ and $M_{N}$. It also extends the compatible parameter space to lower values of $f_{\rm high}$ and $A_{s}^{\rm min}$.
To generalize the discussion above we take $A_{s}>A_{s}^{\rm min}$ and include the dependence on $g^{\prime}$. Note that, although we consider arbitrary values of $g^{\prime}$, the class of models under consideration can admit an elegant UV completion within simple GUT theories for large gauge couplings; $g^{\prime}\gtrsim 10^{-2}$ [58, 59, 60]. Nonetheless, mildly non-standard constructions, such as $E_{6}$ models with $U(1)$ mixing or in string-inspired constructions with non-universal gauge kinetic functions, can accommodate lower values of $g^{\prime}\sim 10^{-3}$ [97, 98, 99]. Thus, although we allow for a small-$g^{\prime}$ regime, the existence of a viable UV completion remains a desirable feature.
In Fig. S2, we select three benchmark values of $f_{\rm high}$ and $A_{s}$ (or, equivalently, $\Omega_{\rm GW,0}(f_{\rm high})$) within the PTA, LISA, and LIGO sensitivity ranges (left to right). Note how, for larger $g^{\prime}$ and for higher $f_{\rm high}$, the mapping of $M_{N}$ onto $f_{\rm high}$ becomes tighter. This supports our main result: for sufficiently large $g^{\prime}$, well-separated leptogenesis scales can be resolved through features in the GW frequency spectrum. In the opposite case, the parameter space becomes more degenerate for smaller values of $g^{\prime}$ and lower $f_{\rm high}$. In this model, however, $f_{\rm high}$ cannot be arbitrarily small: the condition $T_{\rm dec}\gtrsim T_{\rm BBN}$ imposes a stringent constraint on the parameter space, especially for low $g^{\prime}$ (see, e.g., Fig. S1). This requirement also excludes smaller values of $A_{s}\lesssim 10^{-2}$ at the lowest allowed frequencies. An illustrative example is the case where $f_{\rm high}\sim 10^{-8}$ Hz, which lies within the PTA frequency range. In this regime, non-linearities emerge only for large values of $A_{s}$, thereby favoring only large-amplitude induced GW signals, interestingly consistent with those reported by the PTAs [14, 15, 16, 17, 18]. We nevertheless note that, although the PTA frequency range is compatible with relatively small values of $g^{\prime}$, the corresponding leptogenesis scale inferred from the signal spans over two orders of magnitude. Despite being the worst-case resolution, the ability to bound the leptogenesis scale within two orders of magnitude is itself a notable achievement.
Figure S1: Allowed parameter space for smaller $g^{\prime}$ values, specifically $g^{\prime}=10^{-3}$ (left) $g^{\prime}=10^{-4}$ (right). The rest of the description remains the same as in Fig. 1, barring the appearance of $T_{\rm dec}\gtrsim T_{\rm BBN}$ and $v_{\Phi}\lesssim v_{\Phi}^{\rm max}$ constraints in both the plots. In both panels of Fig. S1, the minimum values of $A_{s}^{\rm min}$ appear similar due to the constraint imposed in the parameter scan, namely $A_{s}\in[\max(10^{-9},A_{s}^{\rm min}),\,10^{-2}]$.
Finally, Fig. S3 illustrates how the allowed parameter space reduces by lowering the amplitude of the GW signal. This occurs because a lower GW amplitude excludes a larger region of the $(M_{N},y_{N})$ parameter space via the constraint $A_{s}>A_{s}^{\rm min}(M_{N},f_{N})$. This is evident in the middle and right panels of Fig. S3, which show the shrinking of the allowed parameter space for three benchmark signal amplitudes, with frequencies comparable to those used in Fig. S2. The left panel does not exhibit this trend as clearly, since all three amplitudes must naturally lie within the ballpark of PTA signal strength, as discussed above.
Figure S2: Allowed parameter space in $g^{\prime}$$M_{N}$ plane with a color gradient representing $y_{N}$, for three benchmark pairs of $(\Omega_{\rm GW},f_{\rm high})$ across different frequency bands. These plots illustrate that, as $g^{\prime}$ increases, the mapping of the leptogenesis scale onto the peak GW frequency becomes increasingly sharp.
Figure S3: Allowed parameter space in $g^{\prime}$$M_{N}$ plane, shown without the $y_{N}$ color gradient for comparison with Fig. S2. These plots examine reduced-amplitude benchmarks leading to the shrinking of the parameter space.
I.2 Theoretical uncertainties
In the previous discussion, we treated the GW spectrum as predictable once the model parameters are given. There is, however, some level of theoretical uncertainty. First, the shape and amplitude of the GW spectrum given by Eq. (6) rely on simulations and may carry associated modeling assumptions, such as the initial scale-invariant Gaussian density fluctuations. Thus, besides the simulation’s numerical uncertainties, the primordial spectrum’s shape and possible primordial non-Gaussianities may also be relevant. A thorough evaluation of these effects is beyond the scope of this work. We note, though, that an enhanced, almost scale-invariant primordial spectrum of fluctuations is possible in some models of ultra-slow-roll [95], U(1) axion [112, 113] and two-field inflation [114] (although backreaction effects could be important [115]). Regarding non-Gaussianities, we do not expect any significant impact: the GWs are sourced predominantly by the average, most massive halos, which enter the Hubble radius shortly before decay [48], see, e.g., Refs. [116, 117, 118].
The leptogenesis model may also introduce additional theoretical ambiguities, stemming from assumptions about the hierarchy in the self-interaction and the gauge coupling, as well as the choice of the renormalization scale. These assumptions are adopted as optimal choices to explore a reasonably ample parameter space while maintaining consistency with stringent model constraints. Note, however, that in the adopted scenario, where the field $\Phi$ undergoes smooth rolling during radiation domination, choosing $\lambda\simeq g^{\prime n}$ with $n=3$ represents the safest and most favorable case. This choice avoids the emergence of significant potential barriers that typically arise for $n>3$, while also ensuring that $\lambda$ remains stable against Coleman-Weinberg-type radiative $\mathcal{O}(g^{\prime 4})$ corrections. On the other hand, for $n<2$ and large gauge coupling, in addition to the stronger model constraints, there is a heightened risk of inducing a secondary phase of inflation, which is disfavored in this context. A quantitative exploration of these theoretical uncertainties becomes particularly relevant though when supplemented by an estimate of numerical uncertainties arising from instrumental noise and astrophysical foregrounds, even in the simplest scenario. Such a comparative analysis would allow one to assess potential overlaps between theoretical and experimental uncertainties. A detailed investigation of this kind will be explored elsewhere. Nevertheless, we emphasize that in the case of strong signals with large signal-to-noise ratios, where experimental uncertainties are expected to be small, theoretical uncertainties may play a dominant role in limiting the precision of parameter reconstruction.
II PBHs and their potential stochastic GW background contributions
In the main text, we noted that a sufficiently large spectrum of primordial fluctuations could trigger formation of PBHs. For completeness, below we present indicative estimates of the mass and fraction of the resulting PBHs, drawing on Refs. [86, 119]. We caution, however, that PBH formation during matter domination is intrinsically complex, with nonlinear effects not yet fully understood (see, e.g., Refs. [120, 121, 122]). Consequently, current estimates should be treated with caution, as they depend sensitively on the underlying assumptions.
The mass of the resulting PBHs is given by the mass inside a Hubble volume when a given fluctuation with wavenumber $k$ enters the Hubble radius [89, 123]. In our setup, this is given by
$M_{\rm PBH}=\gamma M_{\rm dec}\left(\frac{k}{k_{\rm dec}}\right)^{-3}\,$, (S1)
where $\gamma\approx 0.2$, $k_{\rm dec}$ is the wavenumber that enters the Hubble radius at decay, that is $k_{\rm dec}=a(T_{\rm dec})H(T_{\rm dec})$, and $M_{\rm dec}$ is the mass inside the Hubble volume at decay, namely
$M_{\rm dec}=5\times 10^{-10}M_{\odot}\left(\frac{g_{\rho}}{106.75}\right)^{-1/2}\left(\frac{T_{\rm dec}}{10^{4}{\rm GeV}}\right)^{-2}\,.$ (S2)
For a scale-invariant spectrum, we expect PBH formation from $k=k_{\rm dom}$ to $k=k_{\rm dec}$. As $M_{\rm PBH}(k_{\rm dom})$ may be extremely small, since $k_{\rm dom}/k_{\rm dec}\approx\sqrt{a_{\rm dec}/a_{\rm dom}}\sim(A_{s}^{\rm min})^{-1/4}$, we focus on the maximum PBH mass generated, that is $M_{\rm PBH}^{\rm Max}=M_{\rm dec}$. Then, since our range of decay temperatures is roughly $10\,{\rm MeV}<T_{\rm dec}<10^{7}\,{\rm GeV}$, we find that $M_{\rm PBH}^{\rm Max}(T_{\rm dec}=10^{7}\,{\rm GeV})=10^{-16\,}M_{\odot}$ and $M_{\rm PBH}^{\rm Max}(T_{\rm dec}=10\,{\rm MeV})=300\,M_{\odot}$, where $M_{\odot}\approx 2\times 10^{33}\,{\rm g}$ is a solar mass. Thus, PBHs are always relatively light if they form, except for the lowest decay temperatures. Interestingly, for $10^{-16}\,M_{\odot}<M_{\rm PBH}<10^{-11}\,M_{\odot}$, PBHs could explain all the dark matter [124]. For $10^{-10}\,M_{\odot}<M_{\rm PBH}<100\,M_{\odot}$, the fraction of PBHs as dark matter, $f_{\rm PBH}$, must be below $10^{-3}$$10^{-2}$.
Following Refs. [86, 119], we estimate the fraction of PBHs as dark matter today to be
$f_{\rm PBH}\approx 2\times 10^{13}\,\beta\,\left(\frac{g_{\rho}}{g_{s}}\right)\left(\frac{T_{\rm dec}}{10^{4}{\rm GeV}}\right)\,$, (S3)
where $\beta\approx 5.7\times 10^{-4}\,A_{s}^{5/2}$ for $10^{-4}<A_{s}<0.25$. For $A_{s}<10^{-4}$, angular momentum during collapse becomes important and can prevent collapse [119]. In fact, a more conservative estimate would be to consider $A_{s}\gtrsim 10^{-3}$ to form PBHs, as other effects like non-sphericity could be important [120, 121, 122]. From Eq. (S3), we see that $f_{\rm PBH}\leq 1$ leads to
$A_{s}(f_{\rm PBH}\leq 1)\lesssim 10^{-4}\left(\frac{T_{\rm dec}}{10^{4}{\rm GeV}}\right)^{-2/5}\,.$ (S4)
Interestingly, we see that for $T_{\rm dec}\sim 10\,{\rm MeV}$ we need $10^{-3}<A_{s}\lesssim 2.4\times 10^{-2}$ to have a substantial fraction of PBHs. For $T_{\rm dec}\gtrsim 30\,{\rm GeV}$ we already need $A_{s}\lesssim 10^{-3}$, which goes beyond the conservative lower bound on $A_{s}$ to trust the estimates. Nevertheless, these estimates and the induced GW signal related to Leptogenesis motivate further studies of PBHs within this context, as PBHs could be an additional signature.
Regarding the possible GW background from the unresolved mergers of PBHs, although its amplitude depends on $f_{\rm PBH}$, we know its peak frequency, as it lies at the so-called “ISCO” frequency, namely at $f_{\rm ISCO}\sim 2.2\,{\rm kHz}M_{\odot}/M_{\rm PBH}$. Thus, for the lowest decay temperature, the peak of the GW background lies at $\sim 10\,{\rm Hz}$. Increasing the decay temperature shifts the peak frequency upward. In the leptogenesis scenario, the GW background from unresolved PBHs therefore lies above $10{\rm Hz}$ and thus does not overlap with the induced GW background. Ref. [125], indicates that the resulting GW background attains amplitudes within the reach of future detectors such as ET, given current constraints on $f_{\rm PBH}$. Accordingly, an induced GW signal observed by PTAs could have a corresponding counterpart from unresolved PBH binaries at ET. If the induced signal peaks at higher frequencies, the corresponding PBH GW background is likewise shifted to higher-frequency bands. We emphasize that these estimates represent necessary conditions for PBH formation, though they may not be sufficient.